Solve the equation.
e^2x - 5e^x +4 = 0

Answer by stanbon(75887)

e^2x - 5e^x +4 = 0
----------------
Factor:
(e^x-4)(e^x-1) = 0
---
e^x = 4 or e^x = 1
------------------------
x = ln(4) or x = ln(1) = 0
==============================
cheers,
Stan H.
===================